# Soft Token Attacks Cannot Reliably Audit Unlearning in Large Language Models

### v1

@@ 4.3 Eliciting random strings
To further show the power of STAs, we use them to elicit random strings. Unlike natural text, the chance of a random string appearing in the training set is negligible. Also, preceding tokens do not inform the selection of the next token. In the following experiments, we pick characters uniformly at random in the range 33-126 of the ASCII table asciitable.com .
To elicit a random string, we initialize the soft prompt using randomly selected tokens. Unlike in the WHP and TOFU experiments, there is only the soft prompt. We then train the soft prompt using AdamW for up to $3000$ iterations per soft token; using learning rate $0.005$, and $\beta{s}=(0.9,0.999)$.
In Figure 2, we report the longest elicited string for a given number of soft tokens. We repeat the experiment five times – e.g., the first marker implies that for each of the five tested random strings of length $150$, we found an effective soft prompt. We observe that not all initializations and seed configurations succeed, in which case a run needs to be restarted with a different seed. If the loss plateaus around $25\%$ of the iterations, we restart the run. However, no single string was restarted more than ten times. We experimented with learning rate schedulers but they did not improve the search.
Our results show that STAs can be used to elicit completely random strings, thus undermining their application for auditing unlearning. Due to limited computational resources and long run-times, our evaluation is limited to $10$ soft tokens, and $400$-character long random strings. This does not provide a bound on the longest string that can be elicited.
Next, we aim to answer why eliciting strings is possible. Prompt-tuning Lester et al. (2021) is a performance efficient fine-tuning technique in which instead of training all weights, one trains only a soft prompt added to the input. STAs can be viewed as an extreme case of prompt-tuning, where instead of training over many prompts, one trains an attack per each prompt. Hence, an LLM that outputs a completion that it was trained on is an expected behavior. We urge against misinterpreting the results and declaring techniques ineffective.
Figure 2: Numbers of random characters generated for the given soft prompt length. A single soft token can force over $150$ random characters – more than any text in the TOFU benchmark. With $10$ soft tokens, it is possible to generate over $400$ random characters.
@@ 5 Discussion & Conclusion
Auditing with hard prompts. Attacks such as greedy coordinate gradient Zou et al. (2023) optimize the attack prompt in the hard token space instead of the soft token space. Hence, they are weaker at eliciting completions. On one hand, this might make them more suitable for auditing unlearning. On the other hand, due to their computational requirements, they are often used to force only the beginning of a harmful completion (e.g. Sure, here’s how to build a bomb…) with the hope that the LLM follows. It is unclear whether this would be sufficient to produce specific unlearned passages. We see it as an interesting direction for future work.
Unlearning vs jail-breaking. Our findings are applicable to the jail-breaking community as well. Prior work Zhang et al. (2024d) hinted that unlearning and preventing harmful outputs can be viewed as the same task – removing or suppressing particular information. STAs and fine-tuning attacks Hu et al. (2024) are useful tools for evaluating LLMs in powerful threat models. It was shown that fine-tuning on benign data, or data unrelated to the unlearning records (for jail-breaking and unlearning respectively) can restore undesirable behavior Łucki et al. (2024).
Variation in gradient-based learning. Prior work showed that removing training records from the training set, and repeating the training can result in the same final model Thudi et al. (2022) depending on the random seed. Even though a record was part of the training run, its influence might be minimal, making unlearning unnecessary. Similarly, it was shown that SGD has intrinsic privacy guarantees, assuming there exists a group of similar records Hyland and Tople (2022). Thus, algorithmic auditing of unlearning might not be possible, and one would have to rely on verified or attested procedures instead Eisenhofer et al. (2023), regardless of their impact on the model.
Distinguishing learned soft tokens. Even though, in most our results, the number of soft tokens required to elicit a completion is the same, we attempted to distinguish between them. To this end, we take all single-token STAs optimized for TOFU (Table 3) and assign a label $y=\{0,1\}$: $y=0$ for $f_{\emptyset}$, and $y=1$ for $f_{ft}$ and the unlearned models. We then train a binary classifier using $f_{\emptyset}$ and $f_{ft}$. While we are able to overfit it and distinguish between $f_{\emptyset}$ and $f_{ft}$, we were not able to train a model that would generalize to the unlearned models, and decisively assign a class. Our approach is similar to Dataset Inference Maini et al. (2021, 2024d) which showed there can be distributional differences between the models, depending on the data they were trained on. Further investigation into what soft tokens are learned during the audit is an interesting direction for future work.
6 Conclusion
In this work, we show that soft token attacks (STAs) cannot reliably distinguish between base, fine-tuned, and unlearned models. In all cases, the auditor can elicit all unlearned information by appending optimized soft prompts to the base prompt. Additionally, we show that STA with a single soft token can elicit $150$ random characters, and over $400$ with soft tokens.
Our work demonstrates that machine unlearning in LLMs needs better evaluation frameworks. While many unlearning methods can be broken by simple paraphrasing of original prompts, or by fine-tuning on partial unlearned data or even unrelated data, STA misrepresents their efficacy.
@@ 7 Limitations & ethical considerations
Limitations. Due to computational constraints our work is limited to 7-8 billion parameter models. Nevertheless, given that LLMs’ expressive power increases with size Kaplan et al. (2020), our results should hold for larger LLMs. Our evaluation with random strings could be extended to verify if there is a clear and generalizable dependency between the number of soft tokens and the maximum number of generated characters.
Ethical considerations. In this work, we show that an auditor (a user) with white-box access to the model, and sufficient compute can elicit any text from the LLM. While it does require knowing the target completion for a given prompt, it is likely that partial completions might be enough, thus allowing the user to elicit harmful information. This may be particularly dangerous in settings where the user has approximate knowledge of the information that had been scrubbed off the LLM.

### v2

@@ 5.3 Eliciting random strings
The chance of a random string appearing in the training set is negligible, and preceding tokens do not inform the selection of the next token. We construct random strings uniformly at random from the range 33-126 of the ASCII table.
We initialize the soft prompt using randomly selected tokens. In this experiment, there is only $x_{s}$, and no $x_{p}$. We then train the soft prompt using AdamW for up to $3000$ iterations per soft token; using $lr=0.005$, and $\beta{s}=(0.9,0.999)$.
In Figure 2, we report the longest elicited string for a given number of soft tokens. We repeat the experiment five times – e.g., the first marker implies that for each of the five tested random strings of length $150$, we found an effective soft prompt. We observe that not all initializations and seed configurations succeed, in which case a run needs to be restarted with a different seed. If the loss plateaus around $25\%$ of the iterations, we restart the run. However, no single string was restarted more than ten times. Our results show that STAs can be used to elicit completely random strings, thus undermining their application for auditing unlearning.
Next, we aim to answer why eliciting strings is possible. Prompt-tuning Lester et al. (2021) is an efficient fine-tuning technique that trains only a soft prompt added to the input instead of all weights. STAs can be viewed as an extreme case of prompt-tuning, where instead of training over many prompts, one trains an attack per each prompt. Thus, an LLM that outputs a completion that it was trained on is an expected behavior. However, one could argue, in practice, a properly unlearned (or aligned) LLM should never output undesirable text.
Figure 2: A single soft token can force over $150$ random characters. With $10$ soft tokens, it is possible to generate over $400$ random characters.
@@ 6 Discussion & Conclusion
Unlearning vs jailbreaking. Our findings also apply to the jailbreaking community. Prior work hinted that unlearning and preventing harmful outputs can be both viewed as suppressing particular information Zhang et al. (2024d). For instance, it was shown that fine-tuning on benign or unrelated data can restore undesirable behavior Hu et al. (2024); Łucki et al. (2024).
Variation in gradient descent. Prior work showed that retraining with some records removed can result in the same final model depending on the seed Thudi et al. (2022). The influence of the records might be minimal, making unlearning unnecessary. Similarly, it was shown that SGD has intrinsic privacy guarantees, assuming there exists a group of similar records Hyland and Tople (2022). Thus, algorithmic auditing of unlearning might not be possible, and one would have to rely on verified or attested procedures instead Eisenhofer et al. (2023), regardless of their impact on the model.
Distinguishing learned soft tokens. Even though, in most of our results, the number of soft tokens required to elicit a completion is the same, we attempt to distinguish between them. To this end, we take all single-token STAs optimized for TOFU (Table 1(b)) and assign a label $y=\{0,1\}$: $y=0$ for $f_{\emptyset}$, and $y=1$ for $f_{ft}$ and the unlearned models. We then train a binary classifier using $f_{\emptyset}$ and $f_{ft}$. While we are able to overfit it and distinguish between $f_{\emptyset}$ and $f_{ft}$, we were not able to train a model that would generalize to the unlearned models, and decisively assign a class. Our approach is similar to Dataset Inference Maini et al. (2021, 2024d) which showed there can be distributional differences between the models, depending on the data they were trained on. Further investigation into what soft tokens are learned during the audit is an interesting direction for future work.
Conclusion. In this work, we show that soft token attacks (STA) cannot distinguish between base, fine-tuned, and unlearned models: a strong auditor can elicit all unlearned text. Also, we show that STA with a single token can elicit $150$ random characters, and over $400$ with $10$ tokens. Our work shows that machine unlearning in LLMs needs careful assumptions to avoid misleading results.
@@ 7 Limitations & ethical considerations
Limitations. Our experiments are constrained to models with 7–8 billion parameters. Nevertheless, since the expressive power of LLMs generally scales with size Kaplan et al. (2020), we expect our findings to extend to larger models. For efficiency reasons, we restrict our evaluation to at most $10$ soft tokens and random strings of up to $400$ characters, which does not establish an upper bound on the length of strings that can be elicited. Future work could investigate whether black-box optimization methods—such as zeroth-order optimization Chen et al. (2017)—can reproduce the elicitation observed in the white-box setting. In addition, extending our evaluation with random strings may help determine whether there exists a clear and generalizable relationship between the number of soft tokens and the maximum number of generated characters.
Ethical considerations. In this work, we show that an auditor with white-box access and sufficient computational resources can elicit arbitrary text from an LLM. Although this requires knowledge of the target completion, partial completions may suffice, enabling the extraction of harmful information—especially when the auditor has approximate prior knowledge of the removed content.
@@ 8 Acknowledgments
Appendix A Auditing process
Figure 3 gives a complete overview of the auditing procedure, and the difference between $A_{o}$ and $A_{\textsl{STA}}$. In Table 2, we summarize the notation.